OpenStudy (anonymous):

Can someone please help!! Find the general solution of the system of differential equations x'=10x−14y y'=25x−35y (where primes indicate derivatives with respect to t) by using the initial conditions x(0)=A y(0)=B x(t)= y(t)=

OpenStudy (anonymous):

This is what I got so far

OpenStudy (dumbcow):

on the 2nd column of your work, you didn't distribute a negative the x's should cancel leaving x'' +25x' = 0

OpenStudy (anonymous):

so r would then equal +/- 5i and then what would we do?

OpenStudy (anonymous):

and thanks sooo much for replying!

OpenStudy (dumbcow):

wait no, r = 0,-25 r^2 +25r = 0 r(r+25) = 0

OpenStudy (anonymous):

ok but what do we do with those values to get x(t) and y(t)?

OpenStudy (dumbcow):

you can form the general solution for x(t) --> x(t) = c1e^-25t + c2 then you have to repeat this process to find y(t)...substitute to get an equation with y'',y' and y this is a great resource http://tutorial.math.lamar.edu/Classes/DE/RealRoots.aspx

OpenStudy (anonymous):

if x(0)=A then A=c1+c2 right

OpenStudy (dumbcow):

correct, looks like you can't find the constants until after you have y(t)

OpenStudy (anonymous):

but how do we solve for y(t)? we knowvthat y'=25x-35y and y=-x'-10x/14 based on the first equation

OpenStudy (dumbcow):

same way we solved for x(t) take 2nd equation, solve for x differentiate and set equal to 1st equation

OpenStudy (anonymous):

so you end up getting y'=25x-35y x=y'+35y/25 x'=y''+35y'/25 x'=10x-14y y''+35y'/25=10(y'+25y/25)-14y

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

do e plug in the y we found earlier?

OpenStudy (anonymous):

we*

OpenStudy (anonymous):

without adding in the value of y we found earlier i ended up with y''+25y'=0

OpenStudy (dumbcow):

looks good...if you continue you will find you get same answer as before y'' +25y' = 0 --> r =0,-25

OpenStudy (dumbcow):

so now y(t) = k1e^-25t +k2 where k1+k2 = B

OpenStudy (anonymous):

but you then get B=c1+c2

OpenStudy (dumbcow):

i used k just to distinguish the constants for x and y im working on it it seems the hard part is solving for the constants

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

so they want the answer in terms of A and B right?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

do we use determinants?

OpenStudy (dumbcow):

i don't think so. ok by differentiating we can find x'(0) and y'(0) x'(0) = -25c1 y'(0) = -25k1 Now from original equations and that x(0) = A and y(0) = B -25c1 = 10A -14B -25k1 = 25A-35B thus we can now solve for c1 and k1 then find c2,k2 by using fact that c2 = A-c1 k2 = B-k1

OpenStudy (anonymous):

so x(t)=-c2-Ae^-25t+A-c1

OpenStudy (dumbcow):

can't have any c's in the answer, only A and B

OpenStudy (anonymous):

but then what ould it be

OpenStudy (anonymous):

would*

OpenStudy (dumbcow):

x'(0) = 10x(0) -14y(0) -25c1 = 10A-14B now solve for c1 by dividing by -25 do you see it now

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

x(t)=((10A-14B)/-25)e^(-25t)+A-(10A-14B/-25)

OpenStudy (dumbcow):

correct...you can simplify the fractions a bit

OpenStudy (dumbcow):

do same thing with y y'(0) = 25x(0) -35y(0)

OpenStudy (anonymous):

but for x(t) the answer ended up being wrong

OpenStudy (dumbcow):

oh do you have the answer

OpenStudy (anonymous):

no its from an online homework set and it only tells u if yr wrong or right

OpenStudy (dumbcow):

you may have needed to simplify it...did you enter it same as you wrote above?

OpenStudy (anonymous):

yea i did but it states that u dont need to simplify it

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=x%27%28t%29+%3D+10x%28t%29-14y%28t%29%2C+y%27%28t%29%3D25x%28t%29-35y%28t%29%2C+x%280%29%3DA%2C+y%280%29%3DB i checked with wolfram, i am correct look at an example so you know what exact form your answer should be in

OpenStudy (dumbcow):

simplified i get \[x(t) = \frac{14B-10A}{25}e^{-25t} + \frac{35A-14B}{25}\]

OpenStudy (dumbcow):

Note: wolfram puts answer in a complicated form

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

thanks a bunch!

OpenStudy (dumbcow):

your welcome